# Ringdown in Vaidya spacetimes: time-dependent frequencies, Penrose limit and time-domain analyses

### v2

@@ B. Frequency domain analyses
Table 2: The values of $\omega_{\rm Sch}^{\rm Im/Re}$ and $\mathcal{C}=(2\ell+1)\omega_{\rm Sch}^{\rm Im}/\omega_{\rm Sch}^{\rm Re}$ of the fundamental mode in the Schwarzschild spacetime with $\ell=4$.
$\omega_{\rm Sch}^{\rm Im}$ $\omega_{\rm Sch}^{\rm Re}$ $\mathcal{C}=(2\ell+1)\omega_{\rm Sch}^{\rm Im}/\omega_{\rm Sch}^{\rm Re}$
Penrose limit $\frac{1}{6\sqrt{3}M}$ $\frac{2\ell+1}{6\sqrt{3}M}$ 1
massless scalar $0.0963917$ $0.867416$ $1.0001$
electromagnetic $0.0958599$ $0.853095$ $1.0113$
metric perturbation $0.0941640$ $0.809178$ $1.0473$
@@
The purpose of this subsection is to compare the results in the time-domain analyses with the frequency domain analyses listed in Table 1. Therefore, we need to clarify the relation between the master variables ${\rm e}^{-\mathrm{i}\tilde{\Omega}T}\psi^{(s)}(x)$ used in the frequency domain analyses with the master variables $\Psi^{(s)}$ defined in this subsection. In particular, we focus on the characteristic frequencies listed in Table 1, that is, the period and damping rate of the oscillation measured at the radius $x={\rm const}$. It should be noted that $x={\rm const.}$ does not imply the constant areal radius $r=2M(\mathcal{V})x$, and the areal radius increases with time proportionally to $M(\mathcal{V})$. In the time-domain analyses, we may define the oscillation frequency and the damping rate by monitoring the amplitudes of the master variable at each oscillation peak as follows. Let us label each peak of the waveform by an integer $n$, and let $\Psi_{n}$ and $\mathcal{V}_{n}$ denote the amplitude and the time at the $n$-th peak. Then we can extract the real frequency $\omega^{\rm Re}_{n}$, imaginary frequency $\omega^{\rm Im}_{n}$ and those ratio $\mathcal{R}_{n}$ from the following equations:
$\omega_{n}^{\rm Re}$ $=$ $\frac{\pi}{\mathcal{V}_{n}-\mathcal{V}_{n-1}},$ (94)
$\omega_{n}^{\rm Im}$ $=$ $\frac{\ln\left|\Psi^{(s)}_{n}/\Psi^{(s)}_{n-1}\right|}{\mathcal{V}_{n}-\mathcal{V}_{n-1}},$ (95)
$\mathcal{R}_{n}$ $:=$ $\left|\frac{\omega^{\rm Im}_{n}}{\omega^{\rm Re}_{n}}\right|=\frac{\ln\left|\Psi^{(s)}_{n-1}/\Psi^{(s)}_{n}\right|}{\pi}.$ (96)
@@
$S=S(\mathcal{V}_{n-1})+(\omega_{n}^{\rm Re}+\mathrm{i}\omega_{n}^{\rm Im})(\mathcal{V}-\mathcal{V}_{n-1})+\mathcal{O}\left((\mathcal{V}-\mathcal{V}_{n-1})^{2}\right),$ (98)
@@
$\omega^{\rm Re}+\mathrm{i}\omega^{\rm Im}=dS/d{\mathcal{V}}.$ (99)
We solve the master equations together with Eq. (81) in the time domain. The initial condition is set on a $\mathcal{V}=\mathcal{V}_{\rm ini}$ null surface as a function of $r(\mathcal{U},\mathcal{V}_{\rm ini})$ as
$\Psi=\mathcal{A}\exp\left[-\frac{\left(r(\mathcal{U},\mathcal{V}_{\rm ini})-3M_{1}\right)^{2}}{2\sigma_{r}^{2}}\right],$ (100)
where $\mathcal{A}$ and $\sigma_{r}$ is set to $0.7$ and $M_{1}/10$. The waveform measured at $x=60$ for the massless scalar field ($s=0$) and $\ell=4$ is shown in Fig. 2.
Figure 2: The waveform measured at $x=60$ for the massless scalar field ($s=0$) and $\ell=4$.
@@
$\Psi^{(0)}\sim r\Phi\sim x\tilde{\Phi}\sim{\rm e}^{-\mathrm{i}\tilde{\Omega}T}\psi^{(0)}(x)={\rm e}^{-\mathrm{i}\hat{\Omega}T}u(x).$ (101)
@@
$\Psi^{(1)}={\rm e}^{-\mathrm{i}\tilde{\Omega}T}\psi^{(1)}(x)={\rm e}^{-\mathrm{i}\hat{\Omega}T}\psi^{(1)}(x).$ (102)
@@
$\mathrm{d}x$ $=$ $\frac{\tilde{f}(x)}{4M}\mathrm{d}\mathcal{V}-\frac{{\rm e}^{\sigma}}{4M}\mathrm{d}\mathcal{U},$ (103)
$\mathrm{d}T$ $=$ $\frac{1}{4M}\mathrm{d}\mathcal{V}+\frac{{\rm e}^{\sigma}}{4M\tilde{f}(x)}\mathrm{d}\mathcal{U}.$ (104)
Comparing the expressions for the metric perturbations (69) and (87), we find
$4M^{2}{\rm e}^{-\mathrm{i}\tilde{\Omega}T}\left(h_{0}(x)\mathrm{d}T+h_{1}(x)\mathrm{d}x\right)=H_{0}(\mathcal{U},\mathcal{V})\mathrm{d}\mathcal{U}+H_{1}(\mathcal{U},\mathcal{V})\mathrm{d}\mathcal{V}.$ (105)
Then, by using the relations (103) and (104), we find
${\rm e}^{-\mathrm{i}\tilde{\Omega}T}h_{1}=(2M)^{-1}\left(H_{1}-\frac{\tilde{f}(x)}{{\rm e}^{\sigma}}H_{0}\right).$ (106)
Substituting the expressions (90) and (91) and using the transformation of the partial derivatives:
${\rm e}^{-\sigma}\frac{\partial\Psi^{(2)}}{\partial\mathcal{U}}$ $=$ $\frac{1}{4M\tilde{f}}\left(\frac{\partial\Psi^{(2)}}{\partial T}-\tilde{f}\frac{\partial\Psi^{(2)}}{\partial x}\right),$ (107)
$\frac{\partial\Psi^{(2)}}{\partial\mathcal{V}}$ $=$ $\frac{1}{4M}\left(\frac{\partial\Psi^{(2)}}{\partial T}+\tilde{f}\frac{\partial\Psi^{(2)}}{\partial x}\right),$ (108)
@@
${\rm e}^{-\mathrm{i}\tilde{\Omega}T}\psi^{(2)}=(4M)^{-1}\left(1-\frac{1}{x}-2M^{\prime}x+x\tilde{f}(x)\frac{\partial}{\partial x}\right)\Psi^{(2)}\sim\left(1+M^{\prime}\frac{\mathcal{V}}{M_{1}}\right)^{-1}\Psi^{(2)},$ (109)
@@
$\Psi^{(2)}\sim{\rm e}^{-\mathrm{i}\tilde{\Omega}T}\left(1+M^{\prime}\frac{\mathcal{V}}{M_{1}}\right)={\rm e}^{-\mathrm{i}\hat{\Omega}T}.$ (110)
@@
$\Psi^{(s)}\sim\exp\left[-\mathrm{i}\hat{\Omega}T\right]=\left(1+M^{\prime}\frac{\mathcal{V}}{M_{1}}\right)^{{\rm Im}\hat{\Omega}/(2M^{\prime})}\left(1+M^{\prime}\frac{\mathcal{V}}{M_{1}}\right)^{-\mathrm{i}{\rm Re}\hat{\Omega}/(2M^{\prime})}.$ (111)
@@
$\omega\sim\left.\frac{\partial}{\partial\mathcal{V}}\left(\hat{\Omega}T\right)\right|_{x={\rm const.}}=\hat{\Omega}\left.\frac{\partial T}{\partial\mathcal{V}}\right|_{x={\rm const.}}=\frac{\hat{\Omega}}{2M(\mathcal{V})}.$ (112)
@@
$\left.\frac{\partial}{\partial\mathcal{V}}{\rm Re}\left({\rm e}^{-\mathrm{i}\hat{\Omega}T}\right)\right|_{\mathcal{V}=\mathcal{V}_{n}}=0,$ (113)
@@
$\frac{{\rm Re}\hat{\Omega}}{2M^{\prime}}\ln\left|1+M^{\prime}\frac{\mathcal{V}_{n}}{M_{1}}\right|=\pi n+\arg\hat{\Omega}$. (114)
@@
$\omega^{\rm Re}_{n}$ $=$ $\frac{\pi M^{\prime}}{M_{1}}\left(1-\exp\left[-\frac{2\pi M^{\prime}}{{\rm Re}\hat{\Omega}}{\rm arg}\hat{\Omega}\right]\right)^{-1}\exp\left[-\frac{2M^{\prime}}{{\rm Re}\hat{\Omega}}{\rm arg}\hat{\Omega}\right]\exp\left[-\frac{2n\pi M^{\prime}}{{\rm Re}\hat{\Omega}}\right$], (115)
$\mathcal{R}_{n}$ $=$ $\left|\frac{\omega^{\rm Im}_{n}}{\omega^{\rm Re}_{n}}\right|=\frac{\ln\left|\Psi^{(s)}_{n-1}/\Psi^{(s)}_{n}\right|}{\pi}=-\frac{{\rm Im}\hat{\Omega}}{{\rm Re}\hat{\Omega}}$. (116)
In Fig. 3, we show the values of $\omega_{n}^{\rm Re/Im}$ and $\mathcal{R}_{n}$ for the massless scalar field (top), Maxwell field (middle), and metric perturbation (bottom) with $M^{\prime}=\pi/2000$ and $\ell=4$. The amplitude of the damping oscillation drops quickly, so that the numerical accuracy cannot be maintained. Therefore, we plot the results with several different values of the initial time $\mathcal{V}_{\rm ini}$, and focus on the time-domain in which the waveform exhibits an idealized exponentially damping oscillation. We can find that, in the time-domain analyses, calculating the value of $\mathcal{R}_{n}$ along the world line of constant $x$, we obtain similar values to those of the frequency domain analyses and the Penrose limit analyses with the correction factor $\mathcal{C}$ estimated from the Schwarzschild case (see Fig. 3).
Figure 3: The values of $\omega_{n}^{\rm Re/Im}$, $\hat{\Omega}^{\rm Re/Im}/(2M(\mathcal{V}))$ and $\Omega^{\rm Re/Im}_{\rm PL}/(2M(\mathcal{V}))$ are plotted for the massless scalar field (top), Maxwell field (middle) and metric perturbation (bottom) with $M^{\prime}=\pi/2000$ and $\ell=4$.
@@ A. Penrose limit
In this section, we treat the Vaidya spacetime with time-dependent accretion rates. To discuss the Penrose limit, first, we need to obtain the trajectory of the dynamical photon sphere reported in Ref. [35]. For the mass function $M(\mathcal{V})$, following Ref. [35], we consider the following functional form:
$M(\mathcal{V})=\left\{\begin{array}[]{cc}M_{1}&\mathcal{V}\leq\mathcal{V}_{1}\\ M_{1}+\frac{1}{2}(M_{2}-M_{1})\left(1-\cos\left(\frac{\mathcal{V}-\mathcal{V}_{1}}{\mathcal{V}_{2}-\mathcal{V}_{1}}\pi\right)\right)&\mathcal{V}_{1}<\mathcal{V}\leq\mathcal{V}_{2}\\ M_{2}&\mathcal{V}_{2}<\mathcal{V}\end{array}\right.$. (117)
@@
$M^{\prime}_{\rm max}=M^{\prime}((\mathcal{V}_{2}+\mathcal{V}_{1})/2)=\frac{\pi}{2}\frac{M_{2}-M_{1}}{\mathcal{V}_{2}-\mathcal{V}_{1}}.$ (118)
The null geodesic equations can be derived from the following Lagrangian:
$\mathcal{L}=\frac{1}{2}\left[-\left(1-\frac{2M(\mathcal{V})}{r}\right)\left(\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\right)^{2}+2\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\frac{\mathrm{d}r}{\mathrm{d}\tau}+r^{2}\left(\frac{\mathrm{d}\phi}{\mathrm{d}\tau}\right)^{2}\right].$ (119)
@@
$\frac{\partial\mathcal{L}}{\partial\dot{r}}$ $=$ $\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau},$ (120)
$\frac{\partial\mathcal{L}}{\partial\dot{\mathcal{V}}}$ $=$ $-\left(1-\frac{2M(\mathcal{V})}{r}\right)\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}+\frac{\mathrm{d}r}{\mathrm{d}\tau}=:E_{\mathcal{V}},$ (121)
$\frac{\partial\mathcal{L}}{\partial\dot{\phi}}$ $=$ $r^{2}\frac{\mathrm{d}\phi}{\mathrm{d}\tau}=:L,$ (122)
where $L$ is a constant of motion while $E_{\mathcal{V}}$ is not a constant. The equations of motion and constraint are written down as follows:
$\frac{\mathrm{d}^{2}\mathcal{V}}{\mathrm{d}\tau^{2}}$ $=$ $-\frac{M(\mathcal{V})}{r^{2}}\left(\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\right)^{2}+\frac{L^{2}}{r^{3}},$ (123)
$\frac{\mathrm{d}^{2}r}{\mathrm{d}\tau^{2}}$ $=$ $\left(1-\frac{2M(\mathcal{V})}{r}\right)\frac{\mathrm{d}^{2}\mathcal{V}}{\mathrm{d}\tau^{2}}-\frac{M^{\prime}(\mathcal{V})}{r}\left(\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\right)^{2}+\frac{2M(\mathcal{V})}{r^{2}}\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\frac{\mathrm{d}r}{\mathrm{d}\tau},$ (124)
$0$ $=$ $\left(1-\frac{2M(\mathcal{V})}{r}\right)\left(\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\right)^{2}-2\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\frac{\mathrm{d}r}{\mathrm{d}\tau}-\frac{L^{2}}{r^{2}},$ (125)
where $M^{\prime}(\mathcal{V})=\mathrm{d}M(\mathcal{V})/\mathrm{d}\mathcal{V}$. We can reproduce the dynamical photon sphere by following the shooting method described in Ref. [35] for $M_{2}=3M_{1}$, $\mathcal{V}_{1}=0$ and $\mathcal{V}_{2}=100M_{1}$ (Fig. 4). We plot the trajectory of the photon sphere given by a function of the time coordinate $\mathcal{V}$ as $r_{\rm PS}(\mathcal{V})$ in Fig. 4.
Figure 4: The photon sphere trajectory for the Vaidya spacetime with $M_{2}=3M_{1}$, $\mathcal{V}_{1}=0$ and $\mathcal{V}_{2}=100M_{1}$. The null dust is accreting in the shaded region.
@@
$k^{\mu}\nabla_{\mu}E^{\mu}_{(a)}=0.$ (126)
Then the metric components $A_{ab}$ can be calculated by using Eq. (142). We plot $A_{11}$ and $A_{22}$ in Fig. 5 for $M_{2}=3M_{1}$, $\mathcal{V}_{1}=0$ and $\mathcal{V}_{2}=100M_{1}$.
@@
We may consider the adiabatic approximation for this dynamical spacetime of the Penrose limit. Specifically, in the small domain around each time, the geometry can be approximated by the metric (24). Then, we can effectively define $\lambda$ and $\Omega$ from the correspondence (39) as follows:
$\lambda_{\rm ad}:=\frac{\alpha_{\rm ad}E_{\mathcal{V}}}{\beta_{\rm ad}L},$ (127)
$\Omega_{\rm ad}:=\frac{E_{\mathcal{V}}}{L},$ (128)
@@
$\alpha_{\rm ad}$ $=$ $\sqrt{A_{11}},$ (129)
$\beta_{\rm ad}$ $=$ $\sqrt{-A_{22}}.$ (130)

### v3

@@ C. Time-domain analyses
Table 2: The values of $\omega_{\rm Sch}^{\rm Im/Re}$ and $\mathcal{C}=(2\ell+1)|\omega_{\rm Sch}^{\rm Im}/\omega_{\rm Sch}^{\rm Re}|$ of the fundamental mode in the Schwarzschild spacetime with $\ell=4$ with the unit $M=1$.
$\omega_{\rm Sch}^{\rm Re}$ $\omega_{\rm Sch}^{\rm Im}$ $\mathcal{C}$
PL $\frac{2\ell+1}{6\sqrt{3}}=0.866025$ $\frac{1}{6\sqrt{3}}=-0.0962250$ 1
Massless scalar $0.867416$ 0.0963917-0.0963917 $1.0001$
Electromagnetic $0.853095$ 0.0958599-0.0958599 $1.0113$
Metric perturbation $0.809178$ 0.0941640-0.0941640 $1.0473$
@@
The purpose of this subsection is to compare the results in the time-domain analyses with the frequency domain analyses listed in Table 1. Therefore, we need to clarify the relation between the master variables ${\rm e}^{-\mathrm{i}\tilde{\Omega}T}\psi^{(s)}(x)$ used in the frequency domain analyses with the master variables $\Psi^{(s)}$ defined in this subsection. In particular, we focus on the characteristic frequencies listed in Table 1, that is, the period and damping rate of the oscillation measured at the radius $x={\rm const}$. It should be noted that $x={\rm const.}$ does not imply the constant areal radius $r=2M(\mathcal{V})x$, and the areal radius increases with time proportionally to $M(\mathcal{V})$. In the time-domain analyses, we may define the oscillation frequency and the damping rate by monitoring the amplitudes of the master variable at each oscillation peak as follows. Let us label each peak of the waveform by an integer $n$, and let $\Psi_{n}$ and $\mathcal{V}_{n}$ denote the amplitude and the time at the $n$-th peak. Then we can extract the real frequency $\omega^{\rm Re}_{n}$, imaginary frequency $\omega^{\rm Im}_{n}$ and the ratio $\mathcal{R}_{n}$ from the following equations:
$\omega_{n}^{\rm Re}$ $=$ $\frac{\pi}{\mathcal{V}_{n}-\mathcal{V}_{n-1}},$ (3.49)
$\omega_{n}^{\rm Im}$ $=$ $\frac{\ln\left|\Psi^{(s)}_{n}/\Psi^{(s)}_{n-1}\right|}{\mathcal{V}_{n}-\mathcal{V}_{n-1}},$ (3.50)
$\mathcal{R}_{n}$ $:=$ $\left|\frac{\omega^{\rm Im}_{n}}{\omega^{\rm Re}_{n}}\right|=\frac{\ln\left|\Psi^{(s)}_{n-1}/\Psi^{(s)}_{n}\right|}{\pi}.$ (3.51)
@@
$S=S(\mathcal{V}_{n-1})+(\omega_{n}^{\rm Re}+\mathrm{i}\omega_{n}^{\rm Im})(\mathcal{V}-\mathcal{V}_{n-1})+\mathcal{O}\left((\mathcal{V}-\mathcal{V}_{n-1})^{2}\right),$ (3.53)
@@
$\omega^{\rm Re}+\mathrm{i}\omega^{\rm Im}=dS/d{\mathcal{V}}.$ (3.54)
We solve the master equations together with Eq. (3.36) in the time domain. The initial condition is set on a $\mathcal{V}=\mathcal{V}_{\rm ini}$ null surface as a function of $r(\mathcal{U},\mathcal{V}_{\rm ini})$ as
$\Psi=\mathcal{A}\exp\left[-\frac{\left(r(\mathcal{U},\mathcal{V}_{\rm ini})-3M_{1}\right)^{2}}{2\sigma_{r}^{2}}\right],$ (3.55)
where $\mathcal{A}$ and $\sigma_{r}$ are set to $0.7$ and $M_{1}/10$. The waveform measured at $x=60$ for the massless scalar field ($s=0$) and $\ell=4$ with $M^{\prime}=\pi/2000$ is shown in Fig. 2. In Fig. 2, for the Schwarzschild case, we plot the waveform $\propto{\rm e}^{-\mathrm{i}\omega_{\rm Sch}\mathcal{V}}$ with the real and imaginary frequency values listed in Table 2 for $M=M_{1}$. The waveforms $\propto{\rm e}^{-\mathrm{i}\omega_{\rm Sch}\mathcal{V}}$ and $\propto{\rm e}^{-\mathrm{i}\hat{\Omega}T(\mathcal{V})}$ are normalized in such a way that they will become unity at $\mathcal{V}=0$.
Figure 2: The waveform measured at $x=60$ for the massless scalar field ($s=0$) and $\ell=4$ with $M^{\prime}=\pi/2000$. For the Schwarzschild case, we plot the waveform $\propto{\rm e}^{-\mathrm{i}\omega_{\rm Sch}\mathcal{V}}$ with the real and imaginary frequency values listed in Tabel 2 for $M=M_{1}$. The waveforms $\propto{\rm e}^{-\mathrm{i}\omega_{\rm Sch}\mathcal{V}}$ and $\propto{\rm e}^{-\mathrm{i}\hat{\Omega}T(\mathcal{V})}$ are normalized so that they will become unity at $\mathcal{V}=0$.
@@
$\Psi^{(0)}\sim r\Phi\sim x\tilde{\Phi}\sim{\rm e}^{-\mathrm{i}\tilde{\Omega}T}\psi^{(0)}(x)={\rm e}^{-\mathrm{i}\hat{\Omega}T}u(x).$ (3.56)
@@
$\Psi^{(1)}={\rm e}^{-\mathrm{i}\tilde{\Omega}T}\psi^{(1)}(x)={\rm e}^{-\mathrm{i}\hat{\Omega}T}\psi^{(1)}(x).$ (3.57)
@@
$\mathrm{d}x$ $=$ $\frac{\tilde{f}(x)}{4M}\mathrm{d}\mathcal{V}-\frac{{\rm e}^{\sigma}}{4M}\mathrm{d}\mathcal{U},$ (3.58)
$\mathrm{d}T$ $=$ $\frac{1}{4M}\mathrm{d}\mathcal{V}+\frac{{\rm e}^{\sigma}}{4M\tilde{f}(x)}\mathrm{d}\mathcal{U}.$ (3.59)
Comparing the expressions for the metric perturbations (3.24) and (3.42), we find
$4M^{2}{\rm e}^{-\mathrm{i}\tilde{\Omega}T}\left(h_{0}(x)\mathrm{d}T+h_{1}(x)\mathrm{d}x\right)=H_{0}(\mathcal{U},\mathcal{V})\mathrm{d}\mathcal{U}+H_{1}(\mathcal{U},\mathcal{V})\mathrm{d}\mathcal{V}.$ (3.60)
Then, by using the relations (3.58) and (3.59), we find
${\rm e}^{-\mathrm{i}\tilde{\Omega}T}h_{1}=(2M)^{-1}\left(H_{1}-\frac{\tilde{f}(x)}{{\rm e}^{\sigma}}H_{0}\right).$ (3.61)
Substituting the expressions (3.45) and (3.46) and using the transformation of the partial derivatives
${\rm e}^{-\sigma}\frac{\partial\Psi^{(2)}}{\partial\mathcal{U}}$ $=$ $\frac{1}{4M\tilde{f}}\left(\frac{\partial\Psi^{(2)}}{\partial T}-\tilde{f}\frac{\partial\Psi^{(2)}}{\partial x}\right),$ (3.62)
$\frac{\partial\Psi^{(2)}}{\partial\mathcal{V}}$ $=$ $\frac{1}{4M}\left(\frac{\partial\Psi^{(2)}}{\partial T}+\tilde{f}\frac{\partial\Psi^{(2)}}{\partial x}\right),$ (3.63)
@@
${\rm e}^{-\mathrm{i}\tilde{\Omega}T}\psi^{(2)}=(4M)^{-1}\left(1-\frac{1}{x}-2M^{\prime}x+x\tilde{f}(x)\frac{\partial}{\partial x}\right)\Psi^{(2)}\sim\left(1+M^{\prime}\frac{\mathcal{V}}{M_{1}}\right)^{-1}\Psi^{(2)},$ (3.64)
@@
$\Psi^{(2)}\sim{\rm e}^{-\mathrm{i}\tilde{\Omega}T}\left(1+M^{\prime}\frac{\mathcal{V}}{M_{1}}\right)={\rm e}^{-\mathrm{i}\hat{\Omega}T}.$ (3.65)
@@
$\Psi^{(s)}\sim\exp\left[-\mathrm{i}\hat{\Omega}T\right]=\left(1+M^{\prime}\frac{\mathcal{V}}{M_{1}}\right)^{{\rm Im}\hat{\Omega}/(2M^{\prime})}\left(1+M^{\prime}\frac{\mathcal{V}}{M_{1}}\right)^{-\mathrm{i}{\rm Re}\hat{\Omega}/(2M^{\prime})}.$ (3.66)
@@
$\omega\sim\left.\frac{\partial}{\partial\mathcal{V}}\left(\hat{\Omega}T\right)\right|_{x={\rm const.}}=\hat{\Omega}\left.\frac{\partial T}{\partial\mathcal{V}}\right|_{x={\rm const.}}=\frac{\hat{\Omega}}{2M(\mathcal{V})}.$ (3.67)
@@
$\left.\frac{\partial}{\partial\mathcal{V}}{\rm Re}\left({\rm e}^{-\mathrm{i}\hat{\Omega}T}\right)\right|_{\mathcal{V}=\mathcal{V}_{n}}=0,$ (3.68)
@@
$\frac{{\rm Re}\hat{\Omega}}{2M^{\prime}}\ln\left|1+M^{\prime}\frac{\mathcal{V}_{n}}{M_{1}}\right|=\pi n+\arg\hat{\Omega}$. (3.69)
@@
$\omega^{\rm Re}_{n}$ $=$ $\frac{\pi M^{\prime}}{M_{1}}\left(1-\exp\left[-\frac{2\pi M^{\prime}}{{\rm Re}\hat{\Omega}}{\rm arg}\hat{\Omega}\right]\right)^{-1}\exp\left[-\frac{2M^{\prime}}{{\rm Re}\hat{\Omega}}{\rm arg}\hat{\Omega}\right]\exp\left[-\frac{2n\pi M^{\prime}}{{\rm Re}\hat{\Omega}}\right$], (3.70)
$\mathcal{R}_{n}$ $=$ $\left|\frac{\omega^{\rm Im}_{n}}{\omega^{\rm Re}_{n}}\right|=\frac{\ln\left|\Psi^{(s)}_{n-1}/\Psi^{(s)}_{n}\right|}{\pi}=-\frac{{\rm Im}\hat{\Omega}}{{\rm Re}\hat{\Omega}}$. (3.71)
In Fig. 3, we show the values of $\omega_{n}^{\rm Re/Im}$ and $\mathcal{R}_{n}$ for the massless scalar field (top), Maxwell field (middle), and metric perturbation (bottom) with $M^{\prime}=\pi/2000$ and $\ell=4$. The amplitude of the damping oscillation drops quickly, so that the numerical accuracy cannot be maintained. Therefore, we plot the results with several different values of the initial time $\mathcal{V}_{\rm ini}$, and focus on the time-domain in which the waveform exhibits an idealized exponentially damping oscillation. We can find that, in the time-domain analyses, calculating the value of $\mathcal{R}_{n}$ along the world line of constant $x$, we obtain similar values to those of the frequency domain analyses and the PL analyses with the correction factor $\mathcal{C}$ estimated from the Schwarzschild case (see Fig. 3).
Figure 3: The values of $\omega_{n}^{\rm Re/Im}$ (left) and $\mathcal{R}_{n}$ (right) for the massless scalar field (top), Maxwell field (middle), and metric perturbation (bottom) with $M^{\prime}=\pi/2000$ and $\ell=4$. In the left panels, the solid lines and short dashed lines show the values of $\hat{\Omega}^{\rm Re/Im}/(2M(\mathcal{V}))$ and $\Omega^{\rm Re/Im}_{\rm PL}/(2M(\mathcal{V}))$, respectively (see Eq. (3.67) for the relation between $\omega$ and $\hat{\Omega}$). The horizontal dashed line shows the value for the Schwarzschild case with $M=M_{1}$.
@@ A. Penrose limit
In this section, we treat the Vaidya spacetime with time-dependent accretion rates. To discuss the PL, first, we need to obtain the trajectory of the dynamical photon sphere reported in Ref. [35]. For the mass function $M(\mathcal{V})$, following Ref. [35], we consider the following functional form:
$M(\mathcal{V})=\left\{\begin{array}[]{cc}M_{1}&\mathcal{V}\leq\mathcal{V}_{1}\\ M_{1}+\frac{1}{2}(M_{2}-M_{1})\left(1-\cos\left(\frac{\mathcal{V}-\mathcal{V}_{1}}{\mathcal{V}_{2}-\mathcal{V}_{1}}\pi\right)\right)&\mathcal{V}_{1}<\mathcal{V}\leq\mathcal{V}_{2}\\ M_{2}&\mathcal{V}_{2}<\mathcal{V}\end{array}\right.$. (4.1)
@@
$M^{\prime}_{\rm max}=M^{\prime}((\mathcal{V}_{2}+\mathcal{V}_{1})/2)=\frac{\pi}{2}\frac{M_{2}-M_{1}}{\mathcal{V}_{2}-\mathcal{V}_{1}}.$ (4.2)
For the metric (3.1), null geodesic equations on the equatorial plane $\theta=\pi/2$ can be derived from the following Lagrangian:
$\mathcal{L}=\frac{1}{2}g_{\mu\nu}\frac{\mathrm{d}x^{\mu}}{\mathrm{d}\tau}\frac{\mathrm{d}x^{\nu}}{\mathrm{d}\tau}=\frac{1}{2}\left[-\left(1-\frac{2M(\mathcal{V})}{r}\right)\left(\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\right)^{2}+2\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\frac{\mathrm{d}r}{\mathrm{d}\tau}+r^{2}\left(\frac{\mathrm{d}\phi}{\mathrm{d}\tau}\right)^{2}\right].$ (4.3)
@@
$\frac{\partial\mathcal{L}}{\partial\dot{r}}$ $=$ $\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau},$ (4.4)
$\frac{\partial\mathcal{L}}{\partial\dot{\mathcal{V}}}$ $=$ $-\left(1-\frac{2M(\mathcal{V})}{r}\right)\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}+\frac{\mathrm{d}r}{\mathrm{d}\tau}=:E_{\mathcal{V}},$ (4.5)
$\frac{\partial\mathcal{L}}{\partial\dot{\phi}}$ $=$ $r^{2}\frac{\mathrm{d}\phi}{\mathrm{d}\tau}=:L,$ (4.6)
where $L$ is a constant of motion while $E_{\mathcal{V}}$ is not a constant. The equations of motion and constraint are written as follows:
$\frac{\mathrm{d}^{2}\mathcal{V}}{\mathrm{d}\tau^{2}}$ $=$ $-\frac{M(\mathcal{V})}{r^{2}}\left(\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\right)^{2}+\frac{L^{2}}{r^{3}},$ (4.7)
$\frac{\mathrm{d}^{2}r}{\mathrm{d}\tau^{2}}$ $=$ $\left(1-\frac{2M(\mathcal{V})}{r}\right)\frac{\mathrm{d}^{2}\mathcal{V}}{\mathrm{d}\tau^{2}}-\frac{M^{\prime}(\mathcal{V})}{r}\left(\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\right)^{2}+\frac{2M(\mathcal{V})}{r^{2}}\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\frac{\mathrm{d}r}{\mathrm{d}\tau},$ (4.8)
$0$ $=$ $\left(1-\frac{2M(\mathcal{V})}{r}\right)\left(\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\right)^{2}-2\frac{\mathrm{d}\mathcal{V}}{\mathrm{d}\tau}\frac{\mathrm{d}r}{\mathrm{d}\tau}-\frac{L^{2}}{r^{2}},$ (4.9)
where $M^{\prime}(\mathcal{V})=\mathrm{d}M(\mathcal{V})/\mathrm{d}\mathcal{V}$. We can reproduce the dynamical photon sphere by following the shooting method described in Ref. [35] for $M_{2}=3M_{1}$, $\mathcal{V}_{1}=0$, and $\mathcal{V}_{2}=100M_{1}$ (Fig. 4). In this case, the value of $M^{\prime}_{\rm max}$ is given by $\pi/100$. We plot the trajectory of the photon sphere given by a function of the time coordinate $\mathcal{V}$ as $r_{\rm PS}(\mathcal{V})$ in Fig. 4.
Figure 4: The photon sphere trajectory for the Vaidya spacetime with $M_{2}=3M_{1}$, $\mathcal{V}_{1}=0$ and $\mathcal{V}_{2}=100M_{1}$. The null dust is accreting in the shaded region. The value of the maximum accretion rate $M^{\prime}_{\rm max}$ is given by $\pi/100$.
@@
$k^{\mu}\nabla_{\mu}E^{\mu}_{(a)}=0.$ (4.10)
Then the functions $A_{ab}$ contained in the metric components shown in Eq. (A.4) can be calculated by using Eq. (A.9). We plot $A_{11}$ and $A_{22}$ in Fig. 5 for $M_{2}=3M_{1}$, $\mathcal{V}_{1}=0$ and $\mathcal{V}_{2}=100M_{1}$.
@@
We may consider the adiabatic approximation for this dynamical spacetime of the PL. Specifically, in the small domain around each time, the geometry can be approximated by the metric (2.24). Then, we can effectively define $\lambda$ and $\Omega$ from the correspondence (2.39) as follows:
$\lambda_{\rm ad}:=\frac{\alpha_{\rm ad}E_{\mathcal{V}}}{\beta_{\rm ad}L},$ (4.11)
$\Omega_{\rm ad}:=\frac{E_{\mathcal{V}}}{L},$ (4.12)
@@
$\alpha_{\rm ad}$ $=$ $\sqrt{A_{11}},$ (4.13)
$\beta_{\rm ad}$ $=$ $\sqrt{-A_{22}}.$ (4.14)